OpenStudy (anonymous):

How do I set this up? Please help: log x + log(x − 3) = 1

OpenStudy (mertsj):

log a + log b = log ab

OpenStudy (mertsj):

First apply that property. What do you get?

OpenStudy (anonymous):

(log x)(log (x-3))?

OpenStudy (anonymous):

drop the log swing the base and pick up its exponent???

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

I'm sorry. I dont understand

OpenStudy (anonymous):

so would it be 10^x-3=1?

OpenStudy (radar):

log a + log b = log ab log x + log(x-3)=log (x^2-3x)=1

OpenStudy (anonymous):

then do i change into exponential form?

OpenStudy (anonymous):

@Mertsj is this right? @radar : 10=x^2-3x for the next step?

OpenStudy (anonymous):

yes.. that's correct.

OpenStudy (anonymous):

now all you gotta do is solve that quadratic... then check your solution(s) for any extraneous ones...

OpenStudy (anonymous):

do you get +/- 13?

OpenStudy (anonymous):

square root of 13 i mean

OpenStudy (anonymous):

10 = x^2 - 3x 0 = x^2 - 3x - 10 0 = (x - 5)(x + 2) x = 5 or x = -2 now check which one works...

OpenStudy (anonymous):

So i can also solve by a^2 + b^2 = c^2? so its 5

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (maheshmeghwal9):

\[\log_{10}{x}+ \log_{10}{(x-3)} =1.\]\[\implies \log_{10}{(x)(x-3)} =1\]Becoz\[\log_a{M}+\log_a{N} =\log_a{MN} . \]therefore\[x(x-3)=10^1.\]\[x^2-3x-10=0\]\[\implies x^2-5x+2x-10=0\]Now do it on ur ow☺

OpenStudy (maheshmeghwal9):

own*

OpenStudy (anonymous):

I did 100=x^2 +3x^2 then solved and got 5

OpenStudy (maheshmeghwal9):

what r u doing! solve like me in my sol.

OpenStudy (anonymous):

Ok i got 5 and -2

OpenStudy (maheshmeghwal9):

ya u r correct. but only one solution will be acceptable which is 5. -2 will not be acceptable becoz In\[\log_{10}{(x-3)}\]If u put x=-2 then (x-3)<0 which say that log does not exist. So as a summary only 5 will be answer